OpenStudy (amigatour):

I'm still wondering why "n" is called nth term. Can anyone elaborate it for me? I'm working on an Arithmetic Sequence.

ganeshie8 (ganeshie8):

consider the below arithmetic sequence : 11, 12, 13, 14, 15, 16,.......

ganeshie8 (ganeshie8):

1st term = 11 2nd term = 12 3rd term = 13 .... n'th term = Tn

ganeshie8 (ganeshie8):

its just a place holder, it means some n'th term is Tn

OpenStudy (amigatour):

Thank you.

ganeshie8 (ganeshie8):

thats all the question ha ? :)

OpenStudy (amigatour):

For now, yes. I ask every time I get confused.

ganeshie8 (ganeshie8):

good :))

OpenStudy (amigatour):

My textbook is full of explanation and I ended up getting frustrated.

OpenStudy (dan815):

you should pay ganeshi to become ur teacher

OpenStudy (dan815):

ganeshi when i have kids i will send them to u lol.. i need you to teach my kids what prime numbers are from age 1, begin with the sentence your age is not a prime number let us begin

OpenStudy (amigatour):

@ Ganeshie8, I already memorize the formulas for this particular problem, but the n was lurking in my head (What's the nth for?).

ganeshie8 (ganeshie8):

omg ! thats a great way to itroduce prime numbers lol textbook is lengthy and discussing certainly helps :)

ganeshie8 (ganeshie8):

which formula u talking about ? \(\large t_n = a + (n-1)d\) this one ?

OpenStudy (amigatour):

yes hahah

OpenStudy (amigatour):

the sum as well

ganeshie8 (ganeshie8):

okay lets go back to our sequence 11, 12, 13, 14, 15.....

ganeshie8 (ganeshie8):

see if u can u identify below :- a = ? d = ?

OpenStudy (dan815):

n is some arbitrary number and nth term is some arbitrary term in a sequence

OpenStudy (amigatour):

a=11

OpenStudy (amigatour):

and so on

OpenStudy (amigatour):

d=1

OpenStudy (dan815):

1st, 2nd,3rd, 4th,nth,6th,7th in this case nth term is the 5th term in this sequence

OpenStudy (amigatour):

Do you want me to calculate the whole problems?

ganeshie8 (ganeshie8):

^^ above see if that strikes something...

OpenStudy (amigatour):

lol... I was just wondering why n is nth.. however, i should just think x and y.

ganeshie8 (ganeshie8):

lets work an example quick, u wil see it :)

OpenStudy (amigatour):

Dan, thank you.

OpenStudy (dan815):

yah

ganeshie8 (ganeshie8):

11, 12, 13, 14, 15..... in this sequence a = 11, d = 1 you're comfortable so far eh

OpenStudy (amigatour):

After that, I will have to get back to my homework. In a few days, I will take my final exam. :)

OpenStudy (dan815):

good luck.. i hope you get 100%

ganeshie8 (ganeshie8):

it wont take long trust me, u already knw everything :) 1st term = 11 2nd term = 12 3rd term = 14 ... 100th term = ?

OpenStudy (amigatour):

Hang on.. :D

ganeshie8 (ganeshie8):

certainly we can use the \(n\)th term formula :- \(\large t_{n} = a + (n-1)d\) plug in \(n = 100\) it gives u 100th term

OpenStudy (dan815):

i will give u a math lesson okay

ganeshie8 (ganeshie8):

plugin \(n = 2\), it gives u \(2\)nd term u can dumbly use that formula, to get any term in the sequence.

OpenStudy (dan815):

when u do math u must think in math

OpenStudy (amigatour):

110=a100

ganeshie8 (ganeshie8):

Perfect ! next find 2nd term

ganeshie8 (ganeshie8):

simply put \(n = 2\) in the formula

OpenStudy (amigatour):

I prefer to understand a concept completely because I don't want to fail nor making my teacher(s) crazy.

ganeshie8 (ganeshie8):

and confirm if u get 12 as expected or not

OpenStudy (amigatour):

a12=12+(100-1)(1) =111

OpenStudy (amigatour):

The sequence is consistent.

OpenStudy (amigatour):

should i say increases consistently at 1?

ganeshie8 (ganeshie8):

yes, d = 1 since its positive, the terms increase consistently at 1

ganeshie8 (ganeshie8):

do u knw the gauss story ?

OpenStudy (amigatour):

I did..yes. I thought it was hilarious.

OpenStudy (amigatour):

I just read it today. :)

ganeshie8 (ganeshie8):

yup ive read it recently... he was able to do the sums when he was like 5 years old kid :o

OpenStudy (amigatour):

He was awesome.

ganeshie8 (ganeshie8):

and the teachers have no clue how to keep him engaged in the class during leisure hours lol

OpenStudy (amigatour):

His formula reminds me of Excel spreadsheet.

OpenStudy (amigatour):

Lol...

ganeshie8 (ganeshie8):

ahhh why excel spreadsheet lol

OpenStudy (amigatour):

He would get medicated if he lives in our days. :D

ganeshie8 (ganeshie8):

lol that was funny ;)

OpenStudy (amigatour):

Did you notice this